### v1

# Spin-photon entanglement of a single Er3+ ion in the telecom band

Spin-photon entanglement of a single Er3+ ion in the telecom band
Mehmet T. Uysal1, Łukasz Dusanowski1∗, Haitong Xu1∗, Sebastian P. Horvath1, Salim Ourari1,
Robert J. Cava2, Nathalie P. de Leon1, Jeff D. Thompson1
These authors contributed equally to this work.jdthompson@princeton.edu
( 1Department of Electrical and Computer Engineering, Princeton University,
Princeton, NJ, 08544, USA
2Department of Chemistry, Princeton University, Princeton, NJ 08544, USA
Long-distance quantum communication using quantum repeaters is an enabling technology for secure communication, distributed quantum computing and quantum-enhanced sensing and metrology [1, 2]. As a building block of quantum repeaters, spin-photon entanglement has been demonstrated with both atomic [3, 4, 5, 6, 7] and solid-state qubits [8, 9, 10, 11, 12, 13, 14]. However, previously demonstrated qubits with long spin coherence do not directly emit photons into the low-loss telecom band that is needed for long-distance communication. Here, we demonstrate spin-photon entanglement using a single Er3+ ion in a solid-state crystal, integrated into a silicon nanophotonic circuit. Direct emission into the telecom band enables an entanglement rate of 1.48 Hz over 15.6 km of optical fiber, with a fidelity of 73(3)%. This opens the door to large-scale quantum networks based on scalable nanophotonic devices and many spectrally multiplexed Er3+ ions [15].
Individually addressable solid-state defects integrated with nanophotonic devices are particularly attractive for combining a single photon source and quantum memory in a scalable platform [16, 17, 18, 19, 20, 21, 14]. However, a challenge is that the typical energy scale for atomic transitions (400 – 900 nm) is outside the low-loss telecom band in standard optical fibers (1.5 $\mu$m), resulting in significant losses for long distance transmission. Two approaches have been demonstrated to circumvent this challenge: single-photon frequency conversion [9, 22, 23, 24, 25, 26, 27] and generating entangled photon pairs using spontaneous parametric downconversion (SPDC), where one photon is in the telecom band and the other matches the wavelength of an atomic memory [13, 28]. However, these approaches suffer from limited efficiency and added noise, and in the case of SPDC, involve non-deterministic sources. Recently, direct telecom spin-photon entanglement of a single emitter has been demonstrated in an InAs/InP quantum dot [29], but direct entanglement with a long-lived spin suitable for long-distance transmission has not been reported.
The rare earth ion (REI) Er3+ provides a direct spin-photon interface in the telecom band, but exploiting individual REIs as single photon sources and quantum memories is challenging because of their low photon emission rate and correspondingly increased sensitivity to decoherence. Recent work has demonstrated incorporating REIs into nanophotonic cavities to enhance the emission rate up to 1000-fold [30, 31, 32, 33, 34, 35, 36, 37], enabling single-shot readout [38, 39] and manipulating nearby nuclear spin ancillae [40, 41, 42], and remote entanglement of two Yb3+:YVO4 defects [14]. Careful selection and engineering of host materials has also led to improved coherence properties. In particular, Er3+:CaWO4 has demonstrated coherence times of 23 ms in small ensembles [43], and low enough spectral diffusion to enable indistinguishable single photon generation in a Si nanophotonic cavity [44]. However, despite these improvements, maintaining the joint spin and optical coherence during photon emission remains a challenge.
Here, we demonstrate spin-photon entanglement directly in the telecom band, with a single Er3+ ion in CaWO4 integrated into a silicon nanophotonic device. The key advance enabling this demonstration is a spin-photon entanglement protocol capable of refocusing spin decoherence while waiting for a photon to be emitted; this protocol has the additional benefit of mitigating an unanticipated optically-induced spin dephasing mechanism. Through the large Purcell factor of the cavity ($P=342$), the absence of losses from frequency conversion, and the low propagation losses at the emission wavelength of $\lambda=1532.6$ nm, we reach a spin-photon entanglement rate of 1.48 Hz over 15.6 kilometers of optical fiber. This rate improves on previously demonstrated telecom band spin-photon entanglement using frequency converted solid-state sources [9, 22, 27] and is comparable to frequency-converted spin-photon entanglement from trapped ions and single atoms in optical tweezers [23, 24, 26]. The fidelity is $F=0.73(3)$, beyond the classical bound of 0.5 [10]. This opens the door to implementing long-distance quantum networks with Er3+ devices.
Our experimental approach is depicted schematically in Fig. 1a. The nanophotonic device architecture has previously been described in Refs. [44, 45]. Briefly, a hybrid structure is formed by a Si nanophotonic device bonded on the CaWO4 substrate (Fig. 1b). The sample is cooled to $T=500$ mK in a 3He cryostat. Light is extracted from the cavity into a bus waveguide and ultimately an optical fiber though a grating coupler, and microwaves to drive spin transitions are applied using a nearby antenna [45]. Er3+ ions are implanted $\sim$10 nm below the surface of the CaWO4 crystal, within the evanescent field of the Si cavity, resulting in a strong enhancement of the emission rate (Fig. 1c). Through careful control of the implantation and annealing conditions, we have observed single ions with spectral diffusion linewidths below 200 kHz and spectral diffusion less than the radiative linewidth on the timescale of photon emission [44]. The ion used in this work has a slightly broader spectral diffusion linewidth of 470 kHz (Extended Data Fig. 1), and a radiative lifetime of $\tau_{r}=18.4$ $\mu$s for the B transition that is resonant with the optical cavity (Fig. 1d), given by the Purcell factor of $P=342$.
In the standard time-bin spin-photon entanglement protocol [46], optical excitation pulses are embedded in a Hahn echo sequence with a free precession period that is long enough to allow the excited state to fully decay between pulses in order to preserve spin-photon correlations. There are two challenges to implementing this directly in our Er3+:CaWO4 devices. First, maintaining coherence for significantly longer than $\tau_{r}$ requires faster dynamical decoupling than a Hahn echo: while the XY-16 coherence time is 200 $\mu$s, the Hahn echo coherence time is only 30 $\mu$s (Fig. 2a), likely limited by fluctuating magnetic field noise from paramagnetic impurities [44]. Second, the optical pulses themselves cause spin decoherence through a previously unreported mechanism, which is not refocused in the standard sequence.
We first investigate spin decoherence caused by optical pulses. We discovered that the insertion of an optical pulse during free precession can reduce the Hahn echo coherence time significantly, to less than 4 $\mu$s (Fig. 2b). The effect does not depend on the optical pulses being resonant with the optical transition of Er3+. The coherence is restored if the optical pulse is applied at the beginning or end of the echo sequence, which suggests a mechanism of spin spectral diffusion induced by the optical pulse (Fig. 2c). Because of its $S_{4}$ site symmetry (Fig. 2d), the spin precession frequency of Er3+:CaWO4 is linearly sensitive to both magnetic and electric fields, where the latter arises from a distortion in the electronic wavefunction that changes the magnetic moment [47, 43]. We disentangle their effects by measuring the induced spectral diffusion, $\sigma_{\omega}$, for different magnetic field orientations in the $ab-$plane, and observe a dependence of $\sigma_{\omega}\propto|\sin\left(2\phi-2\phi_{0}\right)|$ with $\phi_{0}=35^{\circ}(1)$ (Fig. 2e), matching the previously measured electric field induced magnetic moment shift for Er3+:CaWO4 [47]. Therefore, we attribute the optically-induced dephasing to charge noise, and based on the measured scaling with power and pulse duration, conclude that it results from a single-photon absorption process [Extended Data Fig. 2]. We note that previous measurements of spin coherence in Er3+:CaWO4 were performed without optical illumination, and are therefore not sensitive to this effect [44, 48, 43].
[…]
In Fig. 3c, we compare the photon emission timing and spin coherence under optical illumination for both protocols. In a standard Hahn echo time-bin entanglement protocol, the spin coherence decays on the timescale of 20 $\mu$s, which is comparable to $\tau_{r}$. In the modified protocol, the spin coherence is protected beyond $150$ $\mu$s, provided the precise pulse spacing is chosen to avoid resonant interactions with the 183W nuclear spin bath.
We generate and measure spin-photon entanglement using the sequence in Fig. 4a. To measure the photonic state, we send the emitted photons to an unbalanced Mach-Zehnder interferometer (MZI). The time-bin encoded photon, with bins separated by $T=75.5$ $\mu$s, is split between a long arm of 15.6 km and a short arm (attenuated for equivalent loss) and re-combined on a 50:50 beam-splitter (BS) before going to two superconducting nanowire single photon detectors (SNSPD)(Fig. 1a). The detection of a single photon in the early ($[0,T]$) or late ($[2T,3T]$) time bins corresponds to a measurement of the photon in the $Z_{p}$ basis (|$\ket{E}$,|$\ket{L}$), while the detection of a photon in the central time bin, $[T,2T]$, corresponds to a measurement in the superposition basis $\Phi_{p}$, |±$\ket{\pm\phi}=\ket{E}\pm e^{-i\phi}\ket{L})/\sqrt{2}$, where $\phi$ is the relative phase accumulated in the interferometer (Fig. 4b). Instead of stabilizing $\phi$, we track it continuously throughout the experiment using a reference laser sent through the same interferometer, resulting in a uniformly distributed set of measurement bases [Methods S5]. After each entanglement attempt, the spin state is measured only if a photon is detected within the heralding window, using real-time control with a complex programmable logic device (CPLD) [Methods S6]. We compute the entanglement fidelity from the visibility of the entanglement in the X, Y and Z bases (Fig. 4c,d,e), $F=(1+E_{X}+E_{Y}+E_{Z})/4$ [22], where $\langle X_{s}\Phi_{p}\rangle=E_{X}\cos(\phi)$, $\langle Y_{s}\Phi_{p}\rangle=E_{Y}\sin(\phi)$ and $E_{Z}=\langle Z_{s}Z_{p}\rangle$. Here, $\Phi_{p}=\ket{+\phi}\bra{+\phi}-\ket{-\phi}\bra{-\phi}$ |. The experimentally observed visibilities are $\{E_{X},E_{Y},E_{Z}\}=\{0.60(3),0.55(3),0.77(5)\}$ [Methods S8], corresponding to a total fidelity of $F=0.73(3)$. The rate of successful entanglement generation is $R=1.48$ Hz. We note that we only accept photons emitted within the first 2.5 $\mu$s of the collection window, comprising 9.1% of all emitted photons, to minimize sensitivity to errors from fast dephasing of the optical transition [49] [Methods S9].
The measured fidelity is in reasonable agreement with estimated visibilities of $E_{X/Y}=0.63$ and $E_{Z}=0.80$ based on independently measured sources of error [Methods S10]. In particular, we identify significant contributions from spin decoherence and MW pulse errors ($\epsilon_{X/Y}=25$%, $\epsilon_{Z}=14$%), optical decoherence (7%), background counts (5%), initialization errors (3%), and overlap of emission from the early and late time bins (2%).
In this work, we have demonstrated spin-photon entanglement of a single Er3+ ion directly in the telecom band over 15.6 km of optical fiber, enabled by our fast dynamical decoupling protocol that protects spin coherence during photon emission and mitigates optically induced dephasing. The demonstrated entanglement rate is comparable to that obtained with frequency-conversion from single trapped atoms or ions [23, 24, 26] and is higher than that obtained with frequency-converted solid-state sources [9, 22, 27]. The high rate for this distance is achieved by avoiding frequency conversion, and is not limited by the relatively long optical lifetime of the Er3+ ion: after cavity enhancement, the lifetime is shorter than the photon propagation time through a long fiber link such as the one used here. We estimate that the entanglement rate can be increased to 150 Hz with improvements in optical and spin coherence to make use of all emitted photons (×$\times 11$), improvements to the cavity impedance matching and fiber coupling efficiency (×$\times 6$[50], and faster spin reset (×$\times 1.5$) [Methods S11]. The entanglement fidelity is mainly limited by coherence and dark counts, and should improve with the same enhancements.
These results pave the way to future long-distance quantum repeaters based on Er3+ ions interfaced with scalable nanophotonic devices. While the long-term spectral diffusion is still significantly larger than the radiative linewidth, it was recently demonstrated that quasi-static depehasing can be refocused in the context of spin-spin entanglement protocols [14, 51]. A particularly exciting direction is to use multiplexed control of many emitters in the same cavity [15] to store multiple Bell pairs, which can enable higher-rate operation with reduced memory time requirements in a repeater protocol [52], as well as entanglement distillation [53, 54].
Acknowledgements We acknowledge helpful conversations with Shimon Kolkowitz, Rose Ahlefeldt and Adam Turflinger. This work was primarily supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. We also acknowledge support from the DOE Early Career award (DE-SC0020120, for modeling of decoherence mechanisms and spin interactions), as well as AFOSR (FA9550-18-1-0334 and YIP FA9550-18-1-0081), the Eric and Wendy Schmidt Transformative Technology Fund, the Princeton Catalysis Initiative, and DARPA DRINQS (D18AC00015) for establishing the materials spectroscopy pipeline and developing integrated nanophotonic devices. We acknowledge the use of Princeton’s Imaging and Analysis Center, which is partially supported by the PCCM, an NSF MRSEC (DMR-1420541), as well as the Princeton Micro-Nano Fabrication Center.